OpenStudy (jrafferty07):

Please help on Minimum and Maximum question Determine the minimum and maximum values of f(x)=-4(x-6)^2+3 and the graph of g(x)=2cos(2*x-pi)+4

OpenStudy (jameslp1993):

f(x)=-4(x-6)^2+3.. x = 6 and g(x)=2cos(2*x-pi)+4 aplitude: 2. Period: pi. Phase shift pi/2. Hope it helped a little Medel please :)

OpenStudy (jrafferty07):

How do you find the Min and Max for these functions though?

OpenStudy (welshfella):

for the first one the turning point is at (-6,3) and since the coefficient of x^2 is negative this is a maximum maxm of the function = 3. There is no minimum value as its a parabola which opens downwards.

OpenStudy (welshfella):

% turning point is at (6,3)

OpenStudy (jrafferty07):

this is the graph for the second equation

OpenStudy (welshfella):

Yes So you just have to read off the maxm and minm from the graph.

OpenStudy (jrafferty07):

So Maximum would be 6 and Minimum would be 2?

OpenStudy (welshfella):

exactly

OpenStudy (jrafferty07):

and is there an equation to determine the min and max of a function?

OpenStudy (welshfella):

no Not one equation fits all. For a parabola you write the equation in vertex form - as in your question. For a trig function you can either draw the graph or work it out from the amplitude and displacement in some cases.

OpenStudy (welshfella):

Or you can use calculus - have you done any calculus yet?

OpenStudy (welshfella):

finnding derivatives?

OpenStudy (jrafferty07):

I'm in pre-calc this year

OpenStudy (welshfella):

oh ok

OpenStudy (jrafferty07):

But I'm not very confident in finding derivatives

OpenStudy (welshfella):

I#ll see if i can find a good website for you

OpenStudy (jrafferty07):

Thanks

OpenStudy (welshfella):

https://www.mathsisfun.com/calculus/derivatives-rules.html that one summarises it pretty well. When finding maxm and manm you find the derivative of the function and equate it to 0. This will give you the x value of the turning point. One way to find it its a maxm or minm is to find the second derivative and check its sign.

OpenStudy (jrafferty07):

Thanks!

OpenStudy (welshfella):

so for f(x) = -4(x - 6)^2 + 3 derivative = f'(x) = -4*2(x - 6) = 0 -8x + 48 = 0 x = 6 So the turning point of the graph is where x = 6 to find value of f(x) when x = 6 plug x=6 into f(x) - this gives you 3 so turning point when f(x) = 3 to find if this is a maxm or minm find second derivative f'(x) = -8(x - 6) = -8x + 48 Second derivative f" (x) = -8 which is negative so this means f(x) = 3 is a maxm ( negative second derivative would be a minm) and as there is only one value of send derivative there is no minimum

OpenStudy (jrafferty07):

This is my first time using open study so how do I give you a medal

OpenStudy (welshfella):

click on the blue box Best Response

OpenStudy (welshfella):

Make the most of OS - it closes on Jan 31 2017 Im afraid.

OpenStudy (jrafferty07):

why is it closing?

OpenStudy (welshfella):

Basically Money

OpenStudy (jrafferty07):

Oh I believe it

OpenStudy (jrafferty07):

Well thank you for the help!